# Matten: Video Generation with Mamba-Attention

### v1

@@ 2.1 Video Generation
The task of video generation primarily focuses on produce realistic video clips characterized by high-quality visuals and fluid movements. Previous video generation work can be grouped into 3 types. Initially, a number of researchers focused on adapting powerful GAN-based image generation techniques for video creation [17, 18, 19, 20, 21]. Nonetheless, GAN-based methods may lead to problems such as mode collapse, reducing diversity and realism.
@@ 2.2 Mamba
Mamba, a new State-Space Model, has recently gained prominence in deep learning for its universal approximation capabilities and efficient modeling of long sequences, with applications in diverse fields such as medical imaging, image restoration, graphs, NLP, and image generation [34, 35, 36, 37, 38, 39, 40]. Drawing from control systems and leveraging HiPPO initialization [41], these models, like LSSL [11], address long-range dependencies but are limited by computational demands. To overcome this, S4  [42] and other structured state-space models introduce various configurations [43, 44, 9] and mechanisms [10] that have been integrated into larger representation models [45, 46, 47] for tasks in language and speech. Mamba, and its iterations like VisionMamba [12, 13], S4ND [48], and Mamba-ND [49], exhibit a range of computational strategies, from bidirectional SSMs to local convolution and multi-dimensionality considerations. For 3D imaging, T-Mamba [50] tackles the challenges in orthodontic diagnosis due to the powerful ability of Mmaba to handle long-range dependencies. For video understanding, VideoMamba [51] and Video Mamba Suite [52] adapt Mamba to the video domain and address the challenges of local redundancy and global dependencies prevalent in video data. In the domain of diffusion applications using mamba, Zigzag Mamba [53] advances the scalability and efficiency of generating visual content. It tackles the crucial problem of spatial continuity with an innovative scanning approach, incorporates text-conditioning features, and shows enhanced performance across high-resolution image and video datasets.  [54] closely relates to our work, employing the mamba block in the temporal layer of video diffusion. Diverging from previous research focused mainly on local temporal modeling, our method, Matten, is uniquely designed to encompass global temporal dimensions.
@@ 3.2 The Model Variants of Matten
MoCoGAN 124.7 206.6 2886.9 -
VideoGPT 185.9 222.7 2880.6 -
DIGAN 62.5 83.11 1630.2 156.7
StyleGAN-V 47.41 79.52 1431.0 -
PVDM 355.92 75.48 1141.9 540.2
MoStGAN-V 39.70 65.30 1380.3 -
MoCoGAN-HD 111.8 164.1 1729.6 128.1
LVDM - 95.20 372.0 99.0

### v2

@@ 2.1 Video Generation
The task of video generation primarily focuses on producing realistic video clips characterized by high-quality visuals and fluid movements. Previous video generation work can be grouped into 3 types [17, 18]. Initially, a number of researchers focused on adapting powerful GAN-based image generation techniques for video creation [19, 20, 21, 22, 23]. Nonetheless, GAN-based methods may lead to problems such as mode collapse, reducing diversity and realism.
@@ 2.2 Mamba
Mamba, a new State-Space Model, has recently gained prominence in deep learning [36, 37, 38, 39, 40] for its universal approximation capabilities and efficient modeling of long sequences, with applications in diverse fields such as medical imaging, image restoration, graphs, NLP, and image generation [41, 42, 43, 44, 45, 46, 47]. Drawing from control systems and leveraging HiPPO initialization [48], these models, like LSSL [11], address long-range dependencies but are limited by computational demands. To overcome this, S4  [49] and other structured state-space models introduce various configurations [50, 51, 9] and mechanisms [10] that have been integrated into larger representation models [52, 53, 54] for tasks in language and speech. Mamba, and its iterations like VisionMamba [12, 13], S4ND [55], and Mamba-ND [56], exhibit a range of computational strategies, from bidirectional SSMs to local convolution and multi-dimensionality considerations. For 3D imaging, T-Mamba [57] tackles the challenges in orthodontic diagnosis with Mmaba to handle long-range dependencies. For video understanding, VideoMamba [58] and Video Mamba Suite [59] adapt Mamba to the video domain and address the challenges of local redundancy and global dependencies prevalent in video data. In the domain of diffusion applications using mamba, Zigzag Mamba [60] advances the scalability and efficiency of generating visual content. It tackles the crucial problem of spatial continuity with an innovative scanning approach, incorporates text-conditioning features, and shows enhanced performance across high-resolution image and video datasets.  [61] closely relates to our work, employing the mamba block in the temporal layer of video diffusion. Diverging from previous research focused mainly on local temporal modeling, our method, Matten, is uniquely designed to encompass global temporal dimensions.
@@ 3.2 The Model Variants of Matten
MoCoGAN 124.7 206.6 2886.9 - -
VideoGPT 185.9 222.7 2880.6 - -
DIGAN 62.5 83.11 1630.2 156.7 -
StyleGAN-V 47.41 79.52 1431.0 - -
PVDM 355.92 75.48 1141.9 540.2 1077
MoStGAN-V 39.70 65.30 1380.3 - -
MoCoGAN-HD 111.8 164.1 1729.6 128.1 -
LVDM - 95.20 372.0 99.0 5718